OpenStudy (anonymous):

Simplify 7 + 11 – 6 divided by 3 x 4

OpenStudy (anonymous):

7+11-6 = 18-6 =12 ------- 3*4 =12 ....so 12/12 is one...your answer is 1

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

not a problem :) do you know how to do logs?

OpenStudy (anonymous):

no sorry